# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ 5 Discussion and Conclusion
In this paper, we have evaluated LLMs on the complex task of annotating (anti-)solidarity in German parliamentary speeches and interpreted the annotated data to derive new insights into (anti-)solidarity towards migrants in the post-war period as well as recent past. In more detail, we provide both methodological and sociological insights:
Methodological perspective.
We find that instruction-tuned LLMs like Llama-3.3-70B and GPT-4 are effective tools for annotating solidarity frames in parliamentary debates, with open models offering a more cost-effective alternative to closed ones. In our experiments, Llama-3.3-70B performed competitively with GPT-4, especially in few-shot settings. Still, no model consistently outperformed GPT-4, and even GPT-4 did not fully attain the human upper bound. Existing research offers mixed evidence on the reliability of LLMs for annotation tasks: some highlight their limitations in handling subjective or expert-defined taxonomies ziems2024can (109), while others show they can substitute or outperform human coders heseltine2024large (43, 99), or suggest that smaller, fine-tuned models are more reliable in political science applications wang2024selecting (102). In our case, model performance generally improved with model size, but also varied by target group, historical period (with older data proving more difficult), and inference method. These findings highlight the absence of a one-size-fits-all solution: rather than relying on a single model, both model and method selection must be adjusted to the requirements of the specific task.
Moreover, we emphasize the need for validation at every stage of the analysis pipeline. Even keyword search as a first filtering step can introduce biases in the data or can aggregate data that concerns very different sub-populations (such as different groups of migrants in the German post-war period). Regarding the LLM annotation itself, we validated model predictions against a large data set of human annotations (aggregating a year of labor involving multiple annotators), test multiple inference configurations (including zero- and few-shot prompting), and validate long-term trends produced by our two best-performing models (Llama-3.3-70B and GPT-4) against human annotations. We also conducted error analyses to identify where and why models fail to make sure that outputs can be reliably interpreted. This aligns with recent work advocating for human-in-the-loop workflows and partial human validation when using LLMs for annotation tasks, to ensure output quality and interpretability pangakis2024keeping (75, 100).
Despite these efforts, it remains essential to acknowledge model-specific weaknesses when interpreting results. For example, we observed that Llama-3.3-70B occasionally misclassifies anti-solidarity, particularly in instances where criticism of anti-migrant rhetoric is sometimes mistaken for anti-solidarity framing itself. In such cases, models may overlook logical direction or target relevance and assign misleading labels despite otherwise coherent reasoning zhang2025mpvstance (107). One possible reason for this behavior could be the influence of complex political language or long prompt context, which may distract the model from the intended target huang2025breaking (48). Importantly, we could only reveal model errors because we checked the primary data before drawing conclusions.
Social science perspective.
@@
Beyond the methodological contributions, our study also demonstrates how an LLM-based pipeline can support a theoretically informed quantitative analysis through two illustrative case studies: one historical (1949–1957) and one contemporary (2009–2025).
@@
For the early post-war period (1949–1957), we demonstrated how solidarity was shaped and institutionalized in the context of national reconstruction. While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insight for historical and social science research. First, it enables us to move beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity, and shifting rhetorical strategies across different migrant groups. Second, our approach helps integrate migration histories that are often treated separately, such as expellees, DPs, and refugees, into a shared analytical framework, allowing for direct comparison of how solidarity was articulated across groups. This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
@@
For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis. This may be partially explained by solidarity fatigue DeLeur2023 (23) or may be part of a larger right-wing shift Zick2023 (108) on migration issues. While stricter migration stances across Europe have been analyzed before, our study is the first to quantify the shift from solidarity to anti-solidarity in German parliament. Our data also revealed that the shift is most striking in the liberal and conservative parties, motivating further research. This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives aktas2024rise (2), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists bayerlein2021chasing (8). Comparative survey evidence further shows that anti-immigrant attitudes remain a central driver of right-wing populist support across the continent messing2021anti (70).
This type of fine-grained, longitudinal analysis of solidarity would hardly be feasible at such scale using manual annotation alone which, in our project, cost over €29,000 for only 3,500 instances. By contrast, with access to computing resources (e.g., in our case, in-house hardware with NVIDIA A40 GPUs), open-weight models such as Llama-3.3-70B can annotate tens of thousands of instances at minimal additional cost. Even closed models like GPT-4 are far cheaper, with one third of our dataset annotated for about €500. While machine annotation is therefore not ?free?, it is vastly more scalable than human annotation and enables large-scale analyses that would otherwise be financially prohibitive. In this sense, our method bridges ?distant reading? approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded systematic interpretation of parliamentary speech while retaining full corpus coverage.
@@ 6 Limitations and Future Work
Human Label Variation:
Even after months of training, discussion and improvements of the annotation manual, we still observe substantial disagreement between human annotators, especially on the fine-grained level. This indicates that (anti-)solidarity is intrinsically a challenging concept to annotate, especially on parliamentary speeches, and that reasonable disagreement may persist (e.g. whether anti-solidarity should be regarded as group-based or exchange-based when competition between groups around limited resources is highlighted). For this investigation, we used the majority label, but future work should consider methods that explicitly account for human label variation and embrace it plank-2022-problem (76).
Political Bias:
It may be that the LLMs compared in this paper may be subject to political bias as investigated by several prior works motoki2024more (72, 85). In particular, it may be that a left-wing political bias leads to more anti-solidarity being annotated. However, we do not observe a systematic bias toward anti-solidarity in the error analysis (Fig. 6); instead, models often mislabel solidarity and anti-solidarity samples as ?None?. It may also be that the human annotators are biased towards certain labels, but this would need to affect all annotators equally because we do not observe a systematic disagreement towards one label, either. Finally, even if bias were present, this would not explain differences and trends in (anti-)solidarity (which are the main subject of our analysis) but rather the overall level of (anti-)solidarity.
Analysis of (Anti-)solidarity towards women:
In the same way we performed an analysis of (anti-)solidarity trends towards migrants, the (anti-)solidarity towards women could be analyzed. For example, recent work by Diabate2023 (24) suggests that there may be a shift to the left in Germany regarding gender, which would be an opposite trend compared to what we found regarding migration. Due to space and time constraints, such an analysis is not included in this work but appears highly promising for future work.
Further model variations:
We considered a range of models as well as different prompts, and fine-tuning (for Llama models and BERT models) but a vast range of possible model variations remains to be explored, such as ensembles of models, less quantized models, newer models, etc. It may well be that human annotation quality can be fully matched with more refined tools.
Temporal Reliability:
Both model performance (macro F1) and human agreement vary over time, with notable dips in some decades (e.g., the 1960s for Llama-3.3-70B; Fig. 16). This reflects the difficulty of annotating historical language, but overall trend correlations between GPT-4, Llama-3.3-70B, and human annotations remain high. Since our analysis emphasizes broad temporal trends rather than isolated data points, these fluctuations do not undermine the validity of our conclusions.
@@ Appendix A List of Keywords
For Frau (en.: woman) we use Frauen, Frau, Mutter, Mädchen, Mütter, Ehefrau, Müttern, Hausfrauen, Hausfrau, Ehefrauen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Großmutter, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
@@
For Migrant (en.: refugee) we use Flüchtlinge, Ausländer, Flüchtlingen, Zuwanderung, Vertriebenen, Ausländern, Asylbewerber, Migranten, Migration, Heimatvertriebenen, Aussiedler, Einwanderung, Ansiedler, Vertriebene, Zuwanderer, Asylbewerbern, Flüchtling, Heimatvertriebene, Sowjetzonenflüchtlinge, Aussiedlern, Einwanderer, Asylsuchenden, Asylsuchende, Bürgerkriegsflüchtlinge, Zuwanderern, Ansiedlern, Migrantinnen, Vertriebener, Emigranten, Kriegsflüchtlinge, Ausländerinnen, and Immigranten.
When running stability tests (see Subsubsection 4.1.1), we ensure that each sample includes at least 5 keywords, covers at least 10% of the dataset, and spans at least 75% of the timeline, such that enough data is present to create the plots. For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred. The trends are shown in the Fig. 19 in the Appendix.
@@ Appendix B Parties
For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag. For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by poradaautomatische (77).
@@ Appendix C Models Inference and Training Details
All inference runs were conducted on an NVIDIA A40 (80GB) with temperature set to 0.6 and top-p to 0.9. For Mistral-Large-Instruct-2407, we used an NVIDIA A100 to accommodate memory requirements.
@@
Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure 24 for Migrant and in Figure 27 in the Appendix for Woman).
In few-shot settings, one representative example is included for each class label. GPT-4 receives all instructions in a single prompt (see the prompt text in Figure 21 for Migrant and Figure 22 in the Appendix for Woman). Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
@@
For our baseline, we use a BERT-based pipeline with 110M parameters, comprising a high-level category classifier and two subtype models for solidarity and anti-solidarity. All models share a similar architecture, processing inputs with a target (Frau or Migrant) and full text comprising the focus sentence and left and right context. We add a fully connected layer with softmax activation atop the pooled output of the BERT-based models, with 4 output units for each model. To counter class imbalance, minority classes are oversampled to parity with the majority class. We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer AdamW (68) and categorical crossentropy loss.
Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 reimers2019sentence (82). We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions. The model incorporates ReLU activation and dropout (0.1), and maps the embeddings to four and ten target classes for high-level and fine-grained tasks, respectively. To manage class imbalance, we use oversampling. The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
Fine-Tuned Llama Models
We fine-tune Llama-3-8B-It on a single A40 GPU. Llama-3-70B-It is trained across two A40s using PyTorch FSDP and QLoRA [^11] https://www.philschmid.de/fsdp-qlora-llama3. Training spans 3 epochs with a constant learning rate of 2e-4, AdamW optimizer, and max gradient norm of 0.3.

### v2

@@ 5 Discussion and Conclusion
In this paper, we have evaluated LLMs on the complex, theory-driven task of annotating (anti-)solidarity in German parliamentary speeches and assessed whether their outputs support valid downstream social-scientific inference. We then use a validated and statistically corrected annotation pipeline to derive new insights into solidarity and anti-solidarity toward migrants in the postwar period as well as the recent past. In more detail, we provide both methodological and social scientific insights:
Methodological perspective
Our results show that strong instruction-tuned LLMs can reach human agreement in annotating solidarity frames in parliamentary debates, where GPT-5 performs best overall, while open models such as gpt-oss-120B and Llama-3.3-70B provide competitive and more cost-effective alternatives. At the same time, performance varies systematically by target group, historical period, prompting strategy, and model family, with older data and fine-grained labels proving substantially harder. Generally, existing research presents mixed evidence on the reliability of LLMs for annotation tasks: some studies mention limitations in handling subjective or expert-defined taxonomies (Ziems et al., 2024) or find that smaller fine-tuned models are more reliable in political science applications (Wang et al., 2024), while others show that LLMs can substitute for or outperform human coders (Heseltine and Clemm von Hohenberg, 2024; Törnberg, 2024). Our findings are in line with a more cautious interpretation: LLMs are useful annotation tools, but they should not be viewed as neutral coders, nor should their outputs be treated as ?ground truth? data; rather, they are ?measurement tools that require rigorous validation, calibration, and documentation? (Baumann et al., 2025). This idea is reinforced by our error analysis, which shows that even the strongest models in our study make systematic mistakes. For example, Llama-3.3-70B occasionally misclassifies anti-solidarity, mistaking criticism of anti-migrant rhetoric for anti-solidarity itself, and all models show a general tendency to overpredict the None label. More broadly, they underestimate Mixed cases and compress fine-grained subtype distinctions into broader categories, especially group-based subtypes of (anti-)solidarity. These errors become more severe on historical instances, where less explicit forms of solidarity are often missed. More importantly, these errors are not independent across models: they are often correlated, which limits the benefit of ensemble methods. This matters for downstream analysis because small but systematic misclassifications can accumulate into biased temporal trends.
Recent work highlights this problem: high agreement with human annotators can coexist with systematic misclassifications, meaning that benchmark performance alone does not guarantee valid downstream inference (Bavaresco et al., 2025; Gligoric et al., 2025). Even highly accurate LLM-based surrogates can therefore result in misleading substantive conclusions when used directly in downstream analyses, because doing so ignores measurement error between gold-standard and surrogate labels (Egami et al., 2023). Consistent with this, our study finds that DSL-corrected soft-label outputs best recover human-derived trends, especially in fine-grained categories. This suggests that large-scale LLM annotation, combined with a limited set of expert annotations (around 2k in our case), is a viable strategy for downstream inference. The benefits are greatest when soft labels are aggregated across multiple strong models and corrected with DSL, rather than treated as hard predictions from a single model. Recent benchmarking further suggests that DSL performs especially well in low-sample settings, often outperforming alternative correction methods (de Pieuchon et al., 2025). From a resource-allocation perspective, this pipeline combines expensive expert labels with low-cost model labels in a way that improves inference. In our project, expert annotation cost over €29,000 for 3,500 instances, whereas model annotation scaled much more cheaply, whether via open-weight models such as gpt-oss-120B and Llama-3.3-70B on in-house NVIDIA A40 GPUs or via GPT-4, which annotated one third of the dataset for about €500. A smaller set of costly human labels provides the reference basis for validation and correction, while a much larger set of low-cost model labels extends coverage to corpus scale.
Social science perspective
@@
Beyond methodological contributions, our study shows how an LLM-based pipeline can support theoretically informed quantitative analysis through two case studies: one historical (1949–1957) and one contemporary (2009–2025).
@@
For the early postwar period (1949–1957), we showed how solidarity was shaped and institutionalized in national reconstruction. While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insights. First, it moves beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity and shifting rhetorical strategies across migrant groups. Second, our approach integrates migration histories that are often treated separately (expellees, displaced persons, and refugees) into a shared analytical framework, enabling direct comparison of how solidarity was articulated. This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
@@
For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis. This may reflect solidarity fatigue (De Leur et al., 2023) or a larger societal right-wing shift (Zick et al., 2023) on migration. While stricter migration stances across Europe have been studied, our study is the first to quantify this shift in German parliamentary debates. Our data also revealed that the shift is not only explained by the right-wing AfD party entering parliament, but it also visible in the conservative CDU/CSU party and the liberal FDP party, motivating further research. This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives (Aktas, 2024), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists (Bayerlein, 2021). Comparative surveys in Europe likewise show anti-immigrant attitudes remain a central driver of right-wing populist support (Messing and Ságvári, 2021).
Taken together, our findings address a broader question in CSS: whether LLMs can support valid, theory-driven analysis of complex political concepts at scale. Our answer is yes, but only under clear conditions. LLMs are not neutral replacements for expert coders, and their outputs cannot be used uncritically as if they were ground-truth labels, particularly for theory-grounded annotation tasks. Yet when combined with human reference annotation, systematic validation, error analysis, and DSL-based correction, they can support serious social-scientific research by revealing patterns at a scale that would otherwise be accessible only through qualitative analysis. In this sense, our pipeline bridges ?distant reading? approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded and systematic interpretation of parliamentary speech while retaining full corpus coverage.
@@ 6 Limitations and Future Work
Human Coded Reference
Our evaluation and DSL correction rely on human annotations as the best available expert-coded reference. At the same time, some degree of disagreement remains, especially at the fine-grained level, even after extensive training and refinement of the annotation manual. This reflects the intrinsic difficulty of interpreting theoretically informed categories in parliamentary speech and the possibility of reasonable disagreement (e.g.  over whether competition over resources is group- or exchange-based anti-solidarity, or whether a text is better read as compassionate or group-based solidarity.). DSL improves correlation with human annotations, but it cannot remove uncertainty where the reference itself remains interpretive.
Limits of Correction with DSL
DSL improves recovery of human-derived trends, especially relative to raw hard-label predictions, but it does not remove all sources of error. Its performance depends on the quality of the surrogate model outputs, the coverage of the human-coded subset, and the sampling design. In settings with rare labels or sparse historical coverage, correction may remain unstable.
Temporal Reliability
Both model performance and human agreement fluctuate over time (e.g.  the 1960s for several models; Figure˜18), reflecting the challenge of historical language. Although our correction framework mitigates part of this problem, historical variation remains a challenge for both annotation and downstream inference. Still, because our focus is on broad temporal trends rather than individual years, these fluctuations are less consequential for our main substantive conclusions.
(Anti-)Solidarity towards Women
Similar to migrants, one could analyze (anti-)solidarity trends toward women. Diabaté et al. (2023) notes a leftward shift in Germany on gender issues, potentially contrasting with our migration findings. Due to space and time constraints, we do not pursue this here, but consider it a promising avenue for future work.
Political Bias
LLMs compared in this paper may be subject to political bias as investigated by several prior works (Motoki et al., 2024; Rettenberger et al., 2025), with left-wing political bias potentially leading to more anti-solidarity being annotated. Yet our error analysis (Figure˜6) shows no systematic skew toward anti-solidarity; models more often mislabel instances as ?None?. Annotator bias towards certain labels is also possible but would need to be uniform across annotators, which we do not observe. Even if present, bias would affect only overall levels of (anti-)solidarity, not the differences and trends central to our analysis.
@@ Appendix A List of Keywords
For Frau (en.: woman), we use Frauen (and Frau), Mütter (and Mutter, Müttern), Großmutter, Hausfrauen (and Hausfrau), Ehefrauen (and Ehefrau), Mädchen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
@@
For Migrant (en.: migrant) we use Flüchtlinge (and Flüchtling, Flüchtlingen), Ausländer (and Ausländern, Ausländerinnen), Asylsuchende (and Asylsuchende, Asylsuchenden), Asylbewerber (and Asylbewerbern), Heimatvertriebene (and Heimatvertriebenen), Vertriebene (and Vertriebenen, Vertriebener), Aussiedler (and Aussiedlern), Einwanderer (and Einwanderung), Ansiedler (and Ansiedlern), Zuwanderer (and Zuwanderern, Zuwanderung), Migranten (and Migrantinnen, Migration), Sowjetzonenflüchtlinge, Bürgerkriegsflüchtlinge, Kriegsflüchtlinge, Emigranten, and Immigranten.
For robustness, we conducted a stability test by randomly sampling 200 subsets of at least 5 keyword groups (grouping inflectional variants) each, ensuring that each instance covered at least 10% of the dataset and spanned at least 75% of the timeline (such that enough data is present to create the plots). For each subset, we re-estimated the high-level DSL trends (as defined in Section˜3.2.2) and compared the resulting time series to the full-sample trend using Pearson’s correlation coefficient. The correlations remain high on average, with mean values of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; the 25% quantiles are 0.89, 0.85, and 0.75, respectively. This indicates that the main aggregate trends are broadly robust to keyword composition, although robustness is weaker for mixed.
For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred. The trends are shown in the Figure˜21 in the Appendix.
@@ Appendix B Parties
For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag. For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by Porada (2019).
@@ Appendix C Models Inference and Training Details
Inference was conducted on mixed CPU and GPU environments. Closed-source models (gpt-4-1106-preview, gpt-5-chat) were accessed via API.
Open-weight models were run locally:
  • gpt-oss-20B on two NVIDIA A40 (80 GB) GPUs in bfloat16 precision;
  • Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it in full precision on a single A40 GPU;
  • Mistral-Large-Instruct-2407 on a rented NVIDIA A100 in 4-bit quantization to accommodate memory requirements;
  • Llama-3.3-70B-Instruct in 4-bit quantization and Qwen2.5-72B-Instruct on two A40 GPUs in 8-bit quantization;
  • and gpt-oss-120B and Llama-4-Scout-17B-16E-Instruct in 4-bit quantized form (Q4_K_M and Q4_K_XL, respectively) on CPU nodes (32 threads, 128 GB RAM) via llama.cpp [^14] https://github.com/ggml-org/llama.cpp.
All generations used temperature =0.6=0.6 and top-p =0.9=0.9.
@@
Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure˜27 for Migrant and in Figure˜30 in the Appendix for Woman).
In few-shot settings, one representative example is included for each class label. GPT-4 receives all instructions in a single prompt (see the prompt text in Figure˜25 for Migrant and Figure˜26 in the Appendix for Woman). Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
@@
For our baseline, we use a BERT-based pipeline with 110M parameters, comprising a high-level category classifier and two subtype models for solidarity and anti-solidarity. All models share a similar architecture, processing inputs with a target (Frau or Migrant) and full text comprising the focus sentence and left and right context. We add a fully connected layer with softmax activation atop the pooled output of the BERT-based models, with 4 output units for each model. To counter class imbalance, minority classes are oversampled to parity with the majority class. We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer (Loshchilov and Hutter, 2017) and categorical crossentropy loss.
Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 (Reimers and Gurevych, 2019). We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions. The model incorporates ReLU activation and dropout (0.1), and maps the embeddings to four and ten target classes for high-level and fine-grained tasks, respectively. To manage class imbalance, we use oversampling. The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
Appendix D Error Correlation
For two models $m_{i}$ and $m_{j}$, let $\mathcal{E}_{i}$ and $\mathcal{E}_{j}$ denote the sets of instances misclassified by each model, and let $\mathcal{E}^{\text{same}}_{ij}$ be the subset where both models predict the same wrong label. We compute two measures of pairwise error overlap, reported in Table˜4.
Error Jaccard
$J_{ij}=\frac{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}{|\mathcal{E}_{i}\cup\mathcal{E}_{j}|}$
This measures how often two models make errors on the same instances relative to their total error sets.
Same-Wrong Rate
$P_{\text{same-wrong}\mid\text{both-wrong}}=\frac{|\mathcal{E}^{\text{same}}_{ij}|}{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}$
This measures the proportion of jointly misclassified instances on which two models predict the same incorrect label. Table˜4 reports the resulting values for all model pairs.
Model 1 Model 2 Error Jaccard P(Same Wrong $\mid$ Both Wrong)
gpt-oss-120B Qwen-2.5 0.637 0.494
Llama-3.3-70B Qwen-2.5 0.610 0.448
Llama-3.3-70B gpt-oss-120B 0.574 0.456
Table 4: Pairwise overlap in model prediction errors. Error Jaccard measures the proportion of shared errors relative to the union of all errors between two models. P(Same Wrong $\mid$ Both Wrong) measures how often two models predict the same wrong label when both are incorrect.
Appendix E Label Distribution Distance Metrics
We compare model and human label distributions (Section˜3.2.1) using Jensen–Shannon Divergence and total variation distance.
Jensen–Shannon Divergence.
Let $p$ and $q$ denote discrete probability distributions over the same label set, and let $m=\frac{1}{2}(p+q)$. The Jensen–Shannon divergence is defined as
$\mathrm{JSD}(p\,\|\,q)=\frac{1}{2}\mathrm{KL}(p\,\|\,m)+\frac{1}{2}\mathrm{KL}(q\,\|\,m),$
where $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence. Lower values indicate more similar label distributions.
Total Variation Distance.
Total variation distance measures absolute differences in probability mass:
$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{i}|p_{i}-q_{i}|.$
It ranges from 0 (identical distributions) to 1 (disjoint support).
Comparison Total Variation Jensen–Shannon Divergence
Model ensemble vs humans 0.49 (0.32) 0.39 (0.22)
Human vs human (avg 10 splits) 0.45 (0.19) 0.35 (0.12)
Randomized model baseline 0.75 (0.59) 0.69 (0.47)
gpt-oss-120B vs humans 0.58 (0.36) 0.48 (0.29)
Qwen-2.5 vs humans 0.59 (0.42) 0.50 (0.34)
Llama-3.3-70B vs humans 0.57 (0.40) 0.48 (0.32)
GPT-5 vs humans 0.57 (0.33) 0.47 (0.25)
Table 5: Distribution distances between human and model label distributions (lower = closer alignment). Values are shown as fine-grained (high-level).